1.When a 15.1 mL sample of a 0.479 M aqueous nitrous acid solution is titrated with a 0.483 M aqueous sodium hydroxide solution, what is the pH after 22.5 mL of sodium hydroxide have been added?

2.A 21.6 mL sample of 0.214 M diethylamine, (C2H5)2NH, is titrated with 0.256 M perchloric acid.
The pH before the addition of any perchloric acid is
Answer #1

☺ Moles of nitrous a aid = 0.479 M X 15-1816 2 = 0.0072329 males ( Wegk 94d Males of 22591 of OH jons = 0·483M X 22:5X10ML =② In the solution only diethylamine is present. So (C2H5NH + H2O 04+ Cettshith Initial o.214M - Equili (0-214->) M - X X So,